# EEG-Language Modeling for Pathology Detection

### v3

@@ Abstract
Multimodal language modeling has enabled breakthroughs for representation learning, yet remains unexplored in the realm of functional brain data for pathology detection. This paper pioneers EEG-language models (ELMs) trained on clinical reports and 15000 EEGs. We propose to combine multimodal alignment in this novel domain with timeseries cropping and text segmentation, enabling an extension based on multiple instance learning to alleviate misalignment between irrelevant EEG or text segments. Our multimodal models significantly improve pathology detection compared to EEG-only models across four evaluations and for the first time enable zero-shot classification as well as retrieval of both neural signals and reports. In sum, these results highlight the potential of ELMs, representing significant progress for clinical applications.
1 Introduction
Medical neuroimaging, including electroencephalography (EEG), has lagged behind other fields in leveraging the significant advancements of deep learning. While EEG sees widespread clinical use for pathology detection, in particular for epilepsy [1, 2] as well as sleep disorders [3], available annotated data is scarce. As the impressive scaling properties of deep learning are now well described [4, 5], self-supervised learning (SSL) is a promising direction by enabling pretraining with unlabeled data and thereby increasing available training sample sizes [6, 7]. Various such approaches have shown initial success when applied to EEG. These include methods relying on data-augmentations [8, 9], the temporal ordering of EEG data [10], as well as masking and reconstruction [11]. However, these are hindered by the difficulty of creating appropriate data augmentations and, especially reconstruction techniques, by low signal-to-noise. Thus, progress in the medical context has lagged, likely further exacerbated by the modality displaying high similarity between pathologies.
Meanwhile, important further progress was made in computer vision by leveraging natural language as a signal during pretraining [12]. Specifically, contrastive approaches which aim to align embeddings of image-text pairs have shown to yield representations powerful for downstream tasks in radiology [13, 14]. Given that success in radiology is also believed to be bottlenecked by the availability of labeled data and the reliance on fine-grained information [13], this joint modeling approach is a particularly interesting and novel application for the challenging problem of medical EEG. Fortunately, this is made possible by the clinical reports of physicians which accompany hospital EEG recordings and contain information about the patient and recording itself [15].
@@ 1 Introduction
However, language-EEG pretraining also entails unique challenges. First, datasets are generally smaller than those used in radiology and especially computer vision. Second, the clinical reports tend to be highly heterogeneous. While previous applications have paired natural and medical images with short captions [12, 13], EEG reports tend to span multiple paragraphs and include information irrelevant to downstream clinical tasks, potentially hindering the pretraining process. Moreover, they do not contain temporal information about when events occurred during the recording.
@@ 2 Related work
  • Self-supervised learning with EEG data. SSL with EEG data has been predominantly applied to emotion recognition [16, 17], motor imagery [18, 19], sleep staging [9, 19], as well as pathology detection. For the latter application, the temporal order of EEG crops was used initially to demonstrate label-efficient representation learning [10]. Augmentation-based contrastive learning, combined with larger EEG encoders trained on multiple datasets, further improved pathology detection [8]. Recent studies have explored the use of transformers [20, 11], with a focus on scaling while adopting tokenization in an attempt to improve the challenge of effective cross-dataset EEG training.
  • Using EEG for pathology detection. While SSL shows good performance for pathology detection, it is particularly in contexts with little annotated data that it performs well. When more labeled data is available, expert-based feature extraction combined with traditional machine learning classifiers are competitive together with supervised deep learning [21, 22, 23, 24, 25]. This trend has also been observed in other EEG applications [26, 27]. This may indicate that inter-rater variability in EEG classification may create a performance ceiling [28, 22], highlighting the importance of improving classification with limited labels.
  • Medical multimodal language modeling. Medical vision-language modeling aims to guide self-supervised pretraining on medical images using textual information in reports, with performance on a variety of downstream tasks benefiting as a result [29, 30, 13]. Due to less available data in the medical domain, using a pretrained, frozen language encoder was found to boost downstream performance while considerably reducing computational cost [31]. Nevertheless, this line of work has focused mainly on the ECG, X-ray, CT images, and structural MRI images [32, 33, 34].
  • Multiple instance learning. MIL has seen only limited exploration for EEG. Initial studies have investigated the framework by casting crops of EEG as instances and training classifiers for emotion recognition [35], motor imagery [36], mental disorders [37], and sleep apnea [37]. Of these, only the latter has relied on deep learning.
@@ 3.1.1 EEG-language pretraining
Here we detail the setup for pretraining ELMs. Whereas vision-language models are typically trained by aligning a 2D image with a short caption [12, 13], EEG-language modeling is confronted with long EEG time series and multi-paragraph medical reports. To overcome this, we employ text segmentation and time series cropping to create multiple non-overlapping samples per modality and subject. Next, we propose sub-unit alignment by pretraining on these cropped samples. In addition to considerably increasing sample size, this enables the extension of successful approaches in vision-language models. We initially describe two strategies for sub-unit alignment. First, EEG and text representations may be projected using neural networks to a new, shared latent space prior to alignment (as in CLIP; [12, 13]), denoted henceforth as ELMe,l. Alternatively, the EEG embeddings may be projected into the output space of the language model (as in M-FLAG by [31]), denoted as ELMl and trained using a bespoke loss function. This approach was found to reduce latent collapse in smaller data settings [31]. Following a description of these models, we will introduce an extension based on MIL.
@@
[…]
@@
As paired medical EEG data and clinical reports are scarce, training the text encoder function $f_{l}$ from scratch is unlikely to be successful. Furthermore, employing an existing language model and finetuning the model during multimodal pretraining can lead to training instability and collapse of the latent space [38, 31]. To prevent resulting information loss, we follow the recommendations by [31] to use a pretrained language model for $f_{l}$ and freeze its weights during training. For ELMl, we adopt their proposed composite loss to learn $f_{e}$ and $g_{e}$:
@@
$\mathcal{L}_{align}$ $=\|\mathbf{e}-\mathbf{l}\|_{2}^{2}=2-2\mathbf{e}^{\top}\mathbf{l}$ (4)
$\mathcal{L}_{orth}$ $=\sum_{j=1}\left(1-\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)_{jj}\right)^{2}$ (5)
$+\sum_{j\neq k}\left(\mathbf{h}_{e}^{\top}\cdot\mathbf{h}_{e}\right)^{2}_{jk}$ (6)
where $\{j,k\}\in\{1,...,\dim\left(\mathbf{h}_{e}\right)\}^{2}$ and $\mathbf{h}_{e}$ denotes a batch of EEG embeddings. Whereas $\mathcal{L}_{align}$ minimizes the difference between $\mathbf{e}$ and $\mathbf{l}$, $\mathcal{L}_{orth}$ promotes independence between latent dimensions of $\mathbf{h}_{e}$. More specifically, the latter is achieved by manipulating the empirical correlation matrix, where the diagonal and off-diagonal elements are pushed to 1 and 0 respectively [31].
Meanwhile, ELMe,l relies on the cosine similarities between normalized EEG and text embeddings, $s^{e2l}_{j,j}=\mathbf{\hat{e}}_{j}^{\top}\mathbf{\hat{l}}_{j}$, and between text and EEG, $s^{l2e}_{j,j}=\mathbf{\hat{l}}_{j}^{\top}\mathbf{\hat{e}}_{j}$, with $j=1,2,3,...,B$ for batch size $B$ [12]. The multimodal contrastive InfoNCE loss uses a temperature hyperparameter $\tau$ (set to 0.3 using a holdout set; Appendix B.4) and is formulated as:
@@
Multiple instance learning. While previous approaches aim to align text and EEG crops uniformly, certain text segments likely describe specific EEG sections more accurately than others. Therefore, we introduce a MIL alignment strategy that builds on ELMe,l and accommodates multiple positive samples, allowing for more nuanced multimodal relationships. Whereas MIL approaches often rely on operations such as max-pooling to focus on single positive samples, we rely on insights from the video-text alignment approach (MIL-NCE) by [39]. For a given text sample $\mathbf{x}_{l}$, we sample multiple positive EEG crops $\mathbf{x}_{e}$ from the paired recording to approximate the $P(e|l)$ distribution, while for an EEG crop, multiple text segments are sampled to model the $P(l|e)$ distribution. We combine these and sample positives for each EEG and text segment respectively to approximate $P(e,l)$ via bidirectional alignment. This approach effectively relaxes the assumption of strong alignment for each individual $\left(\mathbf{x}_{e},\mathbf{x}_{l}\right)$ pair, instead assuming that, on average, positive samples should have higher similarity scores than negative samples. To this end, we extend the InfoNCE loss to multiple instances:
@@
with |$\text{with }|P_{k}|\leq N\ \text{ and }|Q_{k}|\leq M,$ (12)
@@
and where $P_{k}$ and $Q_{k}$ are sets of positive EEG and text segments respectively. $B_{e}$ and $B_{l}$ are the batch sizes for EEG and text respectively, for which we sample up to $N$ EEG and $M$ text segments for $\frac{B}{N}$ subjects. We set $N=32$ and $M=8$ as this covers all samples for a majority of subjects. We normalize using $|Q_{k}|$ or $|P_{k}|$ to account for the varying number of crops across subjects.
@@
Language encoder. For $f_{l}$ we use a transformer model which was pretrained with medical data in a contrastive manner on PubMed search logs (MedCPT; [40]). See Appendix B.3 for a comparison of language models. ELMl adopts the language model’s native hidden dimensionality (768), while for ELMe,l and ELM-MIL we project to a dimensionality of 256.
@@
3.1.2 EEG-only self-supervised learning
We compare the representations learned by EEG-language pretraining to those obtained via EEG-only pretraining using the same pretraining dataset and EEG encoder. We provide further information on the following methods in Appendix B.5: Bootstrap-Your-Own-Latent (BYOL; [41]), Variance-Invariance-Covariance Regularization (VICReg; [42]), Contrast with the World Representation (ContraWR; [9]), Relative Positioning (RP; [10]), Temporal Shuffling (TS; [10]), Contrastive Predictive Coding (CPC; [10]).
3.2 Datasets and evaluation tasks
• TUEG. The Temple University Hospital (TUH) EEG Corpus is the largest available corpus of hospital EEG data with varying montages, channel counts, and sampling frequencies (n=26846 [15]). For most of the dataset, no labels are available beyond patient age and sex. However, many EEG sessions are associated with a natural-language clinical report.
• TUAB. The TUH Abnormal EEG corpus is a subset of TUEG which was manually labeled by clinicians indicating whether the EEG displays pathological abnormalities [43]. This enables the binary classification task of predicting the status of {normal, abnormal} on a recording-level. Following the literature, we use the provided evaluation set as the hold-out test set.
• NMT. We leverage the NMT Scalp EEG Dataset [44] in order to validate our results for recording-level abnormality classification out-of-distribution. The NMT dataset deviates considerably from TUEG. Data was recorded from a South Asian population in Pakistan, using a different EEG recording setup. Furthermore, the NMT participants are considerably younger, feature more males (66.6%), and their EEG recordings are labeled predominantly normal (83.8% in the training set, while the test set is balanced). We use the provided train/test split.
• TUSZ. The TUH Seizure Detection Corpus [45] is a subset of TUEG which has sections labeled to contain either seizure or background activity. We perform binary classification using 5-fold cross validation on the provided train and dev sets (n=6491), while testing on the eval set.
• TUEV. An Events Corpus [15] containing annotated EEG with six classes, of which three are clinical (spike and slow wave, generalized periodic epileptiform discharge, periodic lateralized epileptiform discharge) as well as eye movements, artifacts, and background activity. We only use the provided train set (5-fold CV) due to the test set not including the TUEG subject identifiers, which would have prevented the exclusion of these subjects from the pretraining data. For each 1-second event, we include two seconds of context before and after [11].
@@ 3.2 Datasets and evaluation tasks
3.3 Preprocessing
3.3.1 Text processing
Due to the heterogeneity of the clinical reports, we further test the utility of summarizing the pathological status indicated by the clinical report using a large language model (LLM). Due to the sensitive nature of the clinical reports, we use the Llama-3 8B model [46] locally and instruct for the production of a single-sentence summary of a report, which should include whether the EEG was deemed abnormal and for which reasons (Appendix E.3).
3.3.2 EEG processing
EEG data received minimal preprocessing, with our approach detailed in Appendix C.1. We describe the selection of our pretraining dataset in Appendix C.2, which avoids data leakage by excluding any data of subjects present in any of the evaluation data. The resulting sample sizes are shown in Table 1. To enable fair comparisons between methods, the optimal crop-length for recording-level pathology detection was determined out of {5,10,20,30,60} seconds without data-leakage (Appendix B.2), yielding 20 and 60 second crops for EEG-only and EEG-language modeling respectively. For TUSZ and TUEV evaluations, we additionally pretrain models using 5-second crops, drop subjects which feature in either TUEV or TUSZ-test, and train with lower learning rates (base learning rate of 0.02 for ELM-MIL and 0.1 otherwise). For ELM-MIL we additionally increase the amount of positive EEG crops $N$ to 120.

### v4

# EEG-Language Pretraining for Highly Label-Efficient Clinical Phenotyping

@@ Abstract
Multimodal language modeling has enabled breakthroughs for representation learning, yet remains unexplored in the realm of functional brain data for clinical phenotyping. This paper pioneers EEG-language models (ELMs) trained on clinical reports and 15000 EEGs. We propose to combine multimodal alignment in this novel domain with timeseries cropping and text segmentation, enabling an extension based on multiple instance learning to alleviate misalignment between irrelevant EEG or text segments. Our multimodal models significantly improve over EEG-only models across four clinical evaluations and for the first time enable zero-shot classification as well as retrieval of both neural signals and reports. In sum, these results highlight the potential of ELMs, representing significant progress for clinical applications.
Machine Learning, ICML
Medical neuroimaging, including electroencephalography (EEG), has lagged behind other fields in leveraging the significant advancements of deep learning. While EEG sees widespread clinical use for clinical phenotyping such as pathology detection, in particular for epilepsy (Binnie & Stefan, 1999; Jing et al., 2020) as well as sleep disorders (Malhotra & Avidan, 2013), available annotated data is scarce. As the impressive scaling properties of deep learning are now well described (Kaplan et al., 2020; Smith et al., 2023), self-supervised learning (SSL) is a promising direction by enabling pretraining with unlabeled data and thereby increasing available training sample sizes (Hadsell et al., 2006; Chen et al., 2020). Various such approaches have shown initial success when applied to EEG. These include methods relying on data-augmentations (Mohsenvand et al., 2020; Yang et al., 2021), the temporal ordering of EEG data (Banville et al., 2021), subject identities of EEG crops (Gijsen & Ritter, 2025), as well as masking and reconstruction (Jiang et al., 2024). However, these are hindered by the difficulty of creating appropriate data augmentations and, especially reconstruction techniques, by low signal-to-noise. Thus, progress in the medical context has lagged, likely further exacerbated by the modality displaying high similarity between pathologies.
Meanwhile, important further progress was made in computer vision by leveraging natural language as a signal during pretraining (Radford et al., 2021). Specifically, contrastive approaches which aim to align embeddings of image-text pairs have shown to yield representations powerful for downstream tasks in radiology (Zhang et al., 2022a, 2023). Given that success in radiology is also believed to be bottlenecked by the availability of labeled data and the reliance on fine-grained information (Zhang et al., 2022a), this joint modeling approach is a particularly interesting and novel application for the challenging problem of medical EEG. Fortunately, this is made possible by the clinical reports of physicians which accompany hospital EEG recordings and contain information about the patient and recording itself (Obeid & Picone, 2016).
@@
However, language-EEG pretraining also entails unique challenges. First, datasets are generally smaller than those used in radiology and especially computer vision. Second, the clinical reports tend to be highly heterogeneous. While previous applications have paired natural and medical images with short captions (Radford et al., 2021; Zhang et al., 2022a), EEG reports tend to span multiple paragraphs and include information irrelevant to downstream clinical tasks, potentially hindering the pretraining process. Moreover, they do not contain temporal information about when events occurred during the recording.
@@ 1 Related work
  • Self-supervised learning with EEG data. SSL with EEG data has been predominantly applied to emotion recognition (Zhang et al., 2022b; Wang et al., 2023), motor imagery (Cheng et al., 2020; Rommel et al., 2022), sleep staging (Yang et al., 2021; Rommel et al., 2022), as well as pathology detection. For the latter application, the temporal order of EEG crops was used initially to demonstrate label-efficient representation learning (Banville et al., 2021). Augmentation-based contrastive learning, combined with larger EEG encoders trained on multiple datasets, further improved pathology detection (Mohsenvand et al., 2020). Still, using subject identities as proxy labels during pretraining was found to improve disease detection over augmentation-based methods (Gijsen & Ritter, 2025). Recent studies have explored the use of transformers (Yang et al., 2024; Jiang et al., 2024), with a focus on scaling while adopting tokenization in an attempt to improve the challenge of effective cross-dataset EEG training.
  • Using EEG for pathology detection. While SSL shows good performance for pathology detection, it is particularly in contexts with little annotated data that it performs well. When more labeled data is available, expert-based feature extraction combined with traditional machine learning classifiers are competitive together with supervised deep learning (Roy et al., 2019; Gemein et al., 2020; Western et al., 2021; Kiessner et al., 2023; Darvishi-Bayazi et al., 2024). This trend has also been observed in other EEG applications (Schirrmeister et al., 2017; Lotte et al., 2018). This may indicate that inter-rater variability in EEG classification may create a performance ceiling (Engemann et al., 2018; Gemein et al., 2020), highlighting the importance of improving classification with limited labels.
  • Medical multimodal language modeling. Medical vision-language modeling aims to guide self-supervised pretraining on medical images using textual information in reports, with performance on a variety of downstream tasks benefiting as a result (Huang et al., 2021; Wang et al., 2022; Zhang et al., 2022a). Due to less available data in the medical domain, using a pretrained, frozen language encoder was found to boost downstream performance while considerably reducing computational cost (Liu et al., 2023a). Nevertheless, this line of work has focused mainly on the ECG, X-ray, CT images, and structural MRI images (Chen et al., 2023; Lalam et al., 2023; Liu et al., 2023b). Recent advances outside the medical domain include multi-task strategies, both during pretraining by integrating contrastive learning and self-supervised losses (Tang et al., 2025; Tschannen et al., 2025), as well as finetuning on multiple downstream tasks (Dai et al., 2023; Liu et al., 2024). Further exploration involves moving compute from unimodal encoding to multimodal fusion (Kim et al., 2021).
  • Multiple instance learning. MIL has seen only limited exploration for EEG. Initial studies have investigated the framework by casting crops of EEG as instances and training classifiers for emotion recognition (Caicedo-Acosta et al., 2019), motor imagery (Collazos-Huertas et al., 2020), mental disorders (Sadatnejad et al., 2019), and sleep apnea (Sadatnejad et al., 2019). Of these, only the latter has relied on deep learning.
@@ 2.1.1 EEG-language pretraining
Here we detail the setup for pretraining ELMs. Whereas vision-language models are typically trained by aligning a 2D image with a short caption (Radford et al., 2021; Zhang et al., 2022a), EEG-language modeling is confronted with long EEG time series and multi-paragraph medical reports. To overcome this, we employ text segmentation and time series cropping to create multiple non-overlapping samples per modality and subject. Next, we propose sub-unit alignment by pretraining on these cropped samples. In addition to considerably increasing sample size, this enables the extension of successful approaches in vision-language models. We initially describe two strategies for sub-unit alignment. First, EEG and text representations may be projected using neural networks to a new, shared latent space prior to alignment (as in CLIP; (Radford et al., 2021; Zhang et al., 2022a)), denoted henceforth as ELMe,l. Alternatively, the EEG embeddings may be projected into the output space of the language model (as in M-FLAG by Liu et al. (2023a)), denoted as ELMl and trained using a bespoke loss function. This approach was found to reduce latent collapse in smaller data settings (Liu et al., 2023a). Following a description of these models, we will introduce an extension based on MIL.
@@
$\mathbf{l}$ $=\begin{cases}g_{l}\left(f_{l}\left(\tilde{\mathbf{x}}_{l}\right)\right)&\text{if ELM${}_{e,l}$}\\ f_{l}\left(\tilde{\mathbf{x}}_{l}\right)&\text{if ELM${}_{l}$}\end{cases}$ (2)
@@
As paired medical EEG data and clinical reports are scarce, training the text encoder function $f_{l}$ from scratch is unlikely to be successful. Furthermore, employing an existing language model and finetuning the model during multimodal pretraining can lead to training instability and collapse of the latent space (Jing et al., 2021; Liu et al., 2023a). To prevent resulting information loss, we follow the recommendations by Liu et al. (2023a) to use a pretrained language model for $f_{l}$ and freeze its weights during training. For ELMl, we adopt their proposed composite loss to learn $f_{e}$ and $g_{e}$:
@@
$\mathcal{L}_{align}$ $=\|\mathbf{\hat{e}}-\mathbf{\hat{l}}\|_{2}^{2}=2-2\mathbf{\hat{e}}^{\top}\mathbf{\hat{l}}$ (4)
$\mathcal{L}_{orth}$ $=\sum_{j=1}\left(1-\left(\mathbf{\hat{h}}_{e}^{\top}\cdot\mathbf{\hat{h}}_{e}\right)_{jj}\right)^{2}$ (5)
$+\sum_{j\neq k}\left(\mathbf{\hat{h}}_{e}^{\top}\cdot\mathbf{\hat{h}}_{e}\right)^{2}_{jk}$ (6)
where $\{j,k\}\in\{1,...,\dim\left(\mathbf{\hat{h}}_{e}\right)\}^{2}$, $\mathbf{\hat{h}}_{e}$ denotes a batch of $\ell_{2}$-normalized EEG embeddings prior to projection and $\mathbf{\hat{e}}$ denotes the normalized projected embeddings. Text embeddings $\mathbf{\hat{l}}$ are likewise normalized. Whereas $\mathcal{L}_{align}$ minimizes the difference between $\mathbf{\hat{e}}$ and $\mathbf{\hat{l}}$, $\mathcal{L}_{orth}$ promotes independence between latent dimensions of $\mathbf{\hat{h}}_{e}$. More specifically, the latter is achieved by manipulating the empirical correlation matrix, where the diagonal and off-diagonal elements are pushed to 1 and 0 respectively (Liu et al., 2023a).
Meanwhile, ELMe,l relies on the cosine similarities between normalized EEG and text embeddings, $s^{e2l}_{j,j}=\mathbf{\hat{e}}_{j}^{\top}\mathbf{\hat{l}}_{j}$, and between text and EEG, $s^{l2e}_{j,j}=\mathbf{\hat{l}}_{j}^{\top}\mathbf{\hat{e}}_{j}$, with j=1,2,3,,Bj=1,2,3,...,B for batch size $B$ (Radford et al., 2021). The multimodal contrastive InfoNCE loss uses a temperature hyperparameter $\tau$ and is formulated as:
@@
Multiple instance learning. While previous approaches aim to align text and EEG crops uniformly, certain text segments likely describe specific EEG sections more accurately than others. Therefore, we introduce a MIL alignment strategy that builds on ELMe,l and accommodates multiple positive samples, allowing for more nuanced multimodal relationships. Whereas MIL approaches often rely on operations such as max-pooling to focus on single positive samples, we rely on insights from the video-text alignment approach (MIL-NCE) by Miech et al. (2020). For a given text sample $\mathbf{x}_{l}$, we sample multiple positive EEG crops $\mathbf{x}_{e}$ from the paired recording to approximate the $P(e|l)$ distribution, while for an EEG crop, multiple text segments are sampled to model the $P(l|e)$ distribution. We combine these and sample positives for each EEG and text report respectively to approximate $P(e,l)$ via bidirectional alignment. This approach effectively relaxes the assumption of strong alignment for each individual $\left(\mathbf{x}_{e},\mathbf{x}_{l}\right)$ pair, instead assuming that, on average, positive samples should have higher similarity scores than negative samples. To this end, we extend the InfoNCE loss to multiple instances:
@@
$\text{with }|P_{k}|\leq N\ \text{ and }|Q_{k}|\leq M,$ (12)
@@
and where $P_{k}$ and $Q_{k}$ are sets of positive EEG and text segments respectively. $B_{e}$ and $B_{l}$ are the batch sizes for EEG and text respectively, for which we sample up to $N$ EEG and $M$ text segments for $\frac{B}{N}$ subjects. We normalize using $|Q_{k}|$ or $|P_{k}|$ to account for the varying number of crops across subjects.
2.1.2 EEG-only self-supervised learning
We compare the representations learned by EEG-language pretraining to those obtained via EEG-only pretraining using the same pretraining dataset and EEG encoder. We emphasize that “EEG-only” refers to pretraining without text, while ELMs use text solely during pretraining to guide EEG representation learning. At test time, neither method uses clinical reports, ensuring alignment with standard EEG-based clinical practice. We provide further information on the following methods in Appendix B.5: Bootstrap-Your-Own-Latent (BYOL; (Grill et al., 2020)), Variance-Invariance-Covariance Regularization (VICReg; (Bardes et al., 2021)), Contrast with the World Representation (ContraWR; (Yang et al., 2021)), Relative Positioning (RP; (Banville et al., 2021)), Temporal Shuffling (TS; (Banville et al., 2021)), Contrastive Predictive Coding (CPC; (Banville et al., 2021)).
3 Experimental Setup
3.1 Datasets and evaluation tasks
• TUEG. The Temple University Hospital (TUH) EEG Corpus is the largest available corpus of hospital EEG data with varying montages, channel counts, and sampling frequencies (n=26846 (Obeid & Picone, 2016)). For most of the dataset, no labels are available beyond patient age and sex. However, many EEG sessions are associated with a natural-language clinical report.
• TUAB. The TUH Abnormal EEG corpus is a subset of TUEG which was manually labeled by clinicians indicating whether the EEG displays pathological abnormalities (Lopez et al., 2015). This enables the binary classification task of predicting the status of {normal, abnormal} on a recording-level. Following the literature, we use the provided evaluation set as the hold-out test set.
• NMT. We leverage the NMT Scalp EEG Dataset (Khan et al., 2022) in order to validate our results for recording-level abnormality classification out-of-distribution. The NMT dataset deviates considerably from TUEG. Data was recorded from a South Asian population in Pakistan, using a different EEG recording setup. Furthermore, the NMT participants are considerably younger, feature more males (66.6%), and their EEG recordings are labeled predominantly normal (83.8% in the training set, while the test set is balanced). We use the provided train/test split.
• TUSZ. The TUH Seizure Detection Corpus (Shah et al., 2018) is a subset of TUEG which has sections labeled to contain either seizure or background activity. We perform binary classification using 5-fold cross validation on the provided train and dev sets (n=6491), while testing on the eval set.
• TUEV. An Events Corpus (Obeid & Picone, 2016) containing annotated EEG with six classes, of which three are clinical (spike and slow wave, generalized periodic epileptiform discharge, periodic lateralized epileptiform discharge) as well as eye movements, artifacts, and background activity. We only use the provided train set (5-fold CV) due to the test set not including the TUEG subject identifiers, which would have prevented the exclusion of these subjects from the pretraining data. For each 1-second event, we include two seconds of context before and after (Jiang et al., 2024).
3.2 Preprocessing
3.2.1 Text processing
Due to the heterogeneity of the clinical reports, we further test the utility of summarizing the pathological status indicated by the clinical report using a large language model (LLM). Due to the sensitive nature of the clinical reports, we use the Llama-3 8B model (Meta, 2024) locally and instruct for the production of a single-sentence summary of a report, which should include whether the EEG was deemed abnormal and for which reasons (Appendix E.4).
3.2.2 EEG processing
EEG data received minimal preprocessing, with our approach detailed in Appendix C.1. We describe the selection of our pretraining dataset in Appendix C.2, which avoids data leakage by excluding any data of subjects present in any of the evaluation data. The resulting sample sizes are shown in Table 1. To enable fair comparisons between methods, the optimal crop-length for recording-level pathology detection was determined out of {5,10,20,30,60} seconds without data-leakage (Appendix B.2), yielding 20 and 60 second crops for EEG-only and EEG-language modeling respectively. For TUSZ and TUEV evaluations, we additionally pretrain models using 5-second crops and drop subjects which feature in either TUEV or TUSZ-test.
3.3 Pretraining setup
We set the temperature parameter $\tau$ to 0.3 using evaluation on a holdout set (Appendix B.4). For ELM-MIL, the number of sampled EEG and text crops per subject depend on the targeted downstream context. When using 60 second crops for subject-level prediction we set the number of sampled EEG crops $N=32$ and text crops $M=8$ paragraphs as this covers the available samples for a majority of subjects. For tasks using single 5-second crops, we increase $N$ to 120 and sample single sentences instead of paragraphs with $M=24$ to accommodate the more finegrained nature of these tasks.
@@ 3.3 Pretraining setup
Language encoder. For $f_{l}$ we use a transformer model which was pretrained with medical data in a contrastive manner on PubMed search logs (MedCPT; (Jin et al., 2023)). See Appendix B.3 for a comparison of language models. ELMl adopts the language model’s native hidden dimensionality (768), while for ELMe,l and ELM-MIL we project to a dimensionality of 256.